OpenStudy (anonymous):

You are given a sample containing 0.90 moles of a substance.? a) How many atoms are in the sample if the substance is copper? b) How many atoms are in the sample if the substance is silver? c) How many molecules are present in the sample if the substance is benzene, C6H6?

OpenStudy (anonymous):

anyone?

OpenStudy (aaronq):

\(n=\dfrac{N}{N_A}\) n=moles, N=particles, \(N_A\)=avogadros number

OpenStudy (anonymous):

for number 1 = .90 = 63.546/6.02x10^23?

OpenStudy (anonymous):

And thanks for helping me earlier!

OpenStudy (aaronq):

no, you're solving for N .90 = N/6.02x10^23?

OpenStudy (anonymous):

so 1.495x10^24 is the answer?

OpenStudy (aaronq):

no, that can't be right. you multiplied by a number less than 1 and got a greater value.

OpenStudy (anonymous):

6.69x10^23?? This right?

OpenStudy (aaronq):

not likely you're still getting more than what you started with

OpenStudy (aaronq):

548002000000000000000000 = 5.48*10^23 should b your answer

OpenStudy (anonymous):

hmm, might be my calculator...

OpenStudy (aaronq):

most likely. you should use an online calculator if it's giving you problems i like this one http://web2.0calc.com or wolfram alpha